# Fast wavelet basis search for generic gravitational wave bursts in Pulsar Timing Array data

### v1

@@ IV Conclusions
With this pipeline developed, it is now computationally feasible to search agnostically for any kind of gravitational wave burst signal in the newest PTA datasets, like the NANOGrav 15 yr dataset, the EPTA DR2 dataset, or the PPTA DR3 dataset [24, 37, 38]. QuickBurst will also allow us to scale this analysis to future combined datasets, like the upcoming IPTA DR3 [39]. Future work includes implementing the ability to simultaneously searching for a generic GW burst signal alongside a common correlated RN process, which would take into account the most recent results found in [1, 2, 3, 4]. Future work also includes expanding this approach to include radio-frequency-dependent wavelets, which can be useful for modeling transient changes in the dispersion measure or other chromatic effect (for a review, see [33]).
@@ Appendix A Per pulsar generic GW burst SNR predictions
In the search for a stochastic background, the SNR of the Hellings-Downs correlations scales with the number of pairs in our PTA [40], which motivates using all the available pulsars in the analysis. However, for a generic GW burst, we need a large enough observing time span for any given pulsar to resolve burst features, and pulsars whose noise properties are understood well enough to be subtracted out as accurately as possible. In addition, removing pulsars decreases the computational cost of these searches, giving us more time to resolve the features in the most significant pulsars. In summary, it may be more efficient to remove pulsars with an observation time span less than the maximum width (given by $\tau$) of a GW burst allowed in a GW burst search. These short timed pulsars will therefore be uninformative regarding the morphology of a GW burst, and can bias $\mathrm{SNR}$ values. These pulsars will then contribute minimally to the total GW burst $\mathrm{SNR}$ [41]. For this reason, the maximum width on the GW signal wavelets is set to $\tau=5$ years for analysis displayed in subsubsection III.2.2.

### v2

@@ IV Conclusions
With this pipeline developed, it is now computationally feasible to search agnostically for any kind of gravitational wave burst signal in the newest PTA datasets, like the NANOGrav 15 yr dataset, the EPTA DR2 dataset, or the PPTA DR3 dataset [24, 37, 38]. QuickBurst will also allow us to scale this analysis to future combined datasets, like the upcoming IPTA DR3 [39]. Future work includes implementing the ability to simultaneously searching for a generic GW burst signal alongside a common correlated RN process, which would take into account the most recent results found in [1, 2, 3, 4]. Future work also includes expanding this approach to include radio-frequency-dependent wavelets, which can be useful for modeling transient changes in the dispersion measure or other chromatic effect (for a review, see [33]). Lastly, DMX is a flexible model that can easily fit features that are not in fact DM variations. This means that DMX can make the covariance between the DM and burst models worse than it intrinsically is, especially on short time scales. Future work may include investigating this covariance through modeling DM as a Gaussian process rather than epoch-by-epoch offsets [40, 41].
@@ Appendix A Per pulsar generic GW burst SNR predictions
In the search for a stochastic background, the SNR of the HD correlations scales with the number of pairs in our PTA [42], which motivates using all the available pulsars in the analysis. However, for a generic GW burst, we need a large enough observing time span for any given pulsar to resolve burst features, and pulsars whose noise properties are understood well enough to be subtracted out as accurately as possible. In addition, removing pulsars decreases the computational cost of these searches, giving us more time to resolve the features in the most significant pulsars. In summary, it may be more efficient to remove pulsars with an observation time span less than the maximum width (given by $\tau$) of a GW burst allowed in a GW burst search. These short timed pulsars will therefore be uninformative regarding the morphology of a GW burst, and can bias $\mathrm{SNR}$ values. These pulsars will then contribute minimally to the total GW burst $\mathrm{SNR}$ [43]. For this reason, the maximum width on the GW signal wavelets is set to $\tau=5$ years for analysis displayed in subsubsection III.2.2.
@@
Appendix B HD correlations covariances with generic wavelet modeling
Recently, the first evidence for the stochastic GWB has been found in PTA data [1, 2, 3, 4]. This motivates considering how HD correlations in the common red noise could impact GW burst analysis. In this analysis, we first consider the spectral characterizations of these common signals to see how significant we expect the impact to be.
As seen in the results from various PTAs, the stochastic GWB has it’s highest contributions in the lower frequency bins based on free spectrum analysis, which are frequencies ranging from $\sim 1-10$ nHz. Because these GW burst searches are intended to search for signals that persist for less than the time span of a dataset, it is not expected that the spectral power from the GWB will significantly overlap with the peak spectral power of a GW burst. An example can be seen in Figure 3, as the wavelet model prefers a frequency between $f_{0}\sim 7-9$ nHz, which does not overlap with the highest power contributions of a stochastic GWB with the most recent spectral characterization of the GWB.
Figure 7: Comparison between two datasets consisting of identical GW burst signals and intrinsic pulsar noise, but only differing in the common noise. One dataset contains an HD correlated signal and the other contains a CURN signal, yet both have the same spectral parameters of $A=6\times 10^{-15}$ and $\gamma=13/3$. (a) Distribution of the match statistic from identical analyses on both datasets. (b) Corner plot of spectral parameter recovery between both datasets. The black crosshair indicates the injected parameter values for both datasets, while the dashed lines represent the median recovered vales for each dataset.
To test this, we create a simulated dataset that is identical in construction to the dataset used to analyze the inclusion of CURN (as shown in Figure 5), with the only difference being including a correlated RN process with HD correlations versus CURN, both with the same spectral parameters. We then perform identical searches with QuickBurst on both datasets. In both cases, we allow the common process in both analyses vary, full knowing QuickBurst only has the ability to model both common processes as CURN.
In Figure 7, we see the results of this analysis. In the top panel, we compute the match statistic for 10,000 randomly drawn samples from the chain returned by QuickBurst for each dataset, and plot the distribution for each. For the dataset containing CURN, the median value of the match statistic is $M=0.84\pm 0.05$, and for the median for the dataset containing HD is $M=0.82\pm 0.04$. We see that the match statistic distributions are consistent within $1\sigma$ of each other. The slight decrease in the match statistic distribution for the dataset with an HD GWB is insignificant due to significant variations between noise realizations in simulations. Therefore, this shows a negligible increase in covariance between the inclusion of HD correlations and our wavelet model.
Moreover, in the bottom panel of Figure 7, we see the spectral recovery for the common signals of each dataset. The black crosshair signifies the injected signal parameters for each, and the dashed lines represent the median recovered spectral parameters for each run. For the CURN dataset, the median recovery for the spectral parameters are $A=3.5\times 10^{-15}$ and $\gamma=4.9$. For the HD dataset, the median recovery for the spectral parameters are $A=2.5\times 10^{-15}$ and $\gamma=5.3$. Given that there are minimal changes in the posteriors for the spectral parameters, this implies that the model for HD correlations absorbs little additional power than the CURN model has shown to absorb from the GW burst that is present. While we can’t conclusively say for sure that a GW burst of any morphology won’t have significant covariance with an HD correlated GWB based on this one example, it is promising that adding in these correlations doesn’t produce any issues with our wavelet model. Similar simulations can easily validate this for other waveforms with different time-frequency content in the future.